# Textoshop: Interactions Inspired by Drawing Software to Facilitate Text Editing

### v1

@@ 5.1.2. Identifying changes was sometimes challenging with Textoshop
Some participants noted how a main difference with drawing software was that it is not as immediate to perceive textual changes. P8 commented “in like Figma […] it’s immediately, visually obvious what changed. […] It didn’t feel like I had the same mobility here where I could like immediately understand what the changes and the consequences were”. In fact, some participants mentioned how they would prefer having some time to review the changes before integrating them. P8 added “I really want to see things side-by-side”. P6 explained that they would like a system to track the changes because the current animation of changes “disappeared so quickly before I can notice [the changes]”. In Textoshop, changes can be tracked either by working on copies or using layers, but these comments seem to indicate that, unlike drawing software, this is not enough to help users identify textual changes.
@@ 5.2.1. Textoshop might not be easy to learn for people with no experience with drawing software
5.2.2. The implementation of Textoshop inherits the issues of using large language models to edit text
While the implementation of Textoshop is not core to our contribution, it is important to acknowledge that Textoshop currently heavily relies on OpenAI’s GPT-4o. It is known that large language models have issues in terms of biases, privacy, and plagiarism (Weidinger et al., 2021; Meyer et al., 2023). We also found the model to perform poorly on tasks such as resizing to a specific number of words. In our implementation, we workaround these issues by regenerating multiple versions until one that matches the length is found. But it is often imperfect, which might frustrate users.
@@ 5.2.3. The drawing software analogy might confuse users
The difficulty with adapting interactions from another domain is that users might not have the same mental model for what a feature might do. This difficulty is coupled with the fact many features are unpredictable due to relying on a large language model, which might change the text in unexpected ways. While this did not seem to be an issue in accomplishing tasks during the study, some participants did mention a mismatch between their mental model and the behaviour of the system. For example, P3 mentioned not seeing text layers like stacked sheets but more like “variables or conditions that you’re kind of swapping in and out”. Similarly, P8 expected features and were surprised not to find them “In Figma […] when I click again it goes down a layer rather than staying at the same level”.
@@ 5.3.1. Repurposing other digital drawing features
We only scratched the surface of the drawing features that could be repurposed for text. For example, layers could support different blending modes to either overwrite or track the changes and changing the opacity of the layer could control how the text is integrated (e.g., 100% overwrites, while 50% generates text with a meaning midway between the layer and what it overlaps). Other layer types could include “Adjustment layers” (Andrews, 2006) that apply a transformation to the text (e.g., making the text more positive) while retaining the integrity of the text underneath the layer. Features such as the rotation could be made more powerful by allowing users to specify the centre of rotation (the word that should remain at the same location). Finally, text could be transformed by manipulating histograms that show text properties such as the number of adverbs, the sentiment of the text, or the length of the sentences.
@@ 5.3.2. Leveraging recent research in graphical editing to also benefit Textoshop
The design of drawing software is a well-researched topic. Much of this research could be adapted and implemented in Textoshop. For example, work on reifying (Beaudouin-Lafon and Mackay, 2000) graphical attributes into manipulable objects (Xia et al., 2016) could be adapted so that properties such as tone, sentence length, and sentence rotation become first-class objects that can be easily manipulated, copied, reversed, and removed. Work on better colour pickers and colour theme authoring could similarly power better tone pickers where tones become objects that can be rearranged and interpolated (Shugrina et al., 2019). Similarly, work on using histograms as another editing space could inspire histograms of text properties that can help find and edit the text (Chevalier et al., 2012).
@@ 5.3.3. Applying the drawing software analogy to other domains
Analogies used to design software often stay within the same domain: the digital canvas for drawing software and the blank sheet for text editors. Through this work, we show the potential of using analogies from different domains. We believe this strategy could inform the design of software from all domains. For example, code editors could also benefit from layers to structure the code. This would allow an organization with a finer granularity than files, and layers could be hidden to remove the code responsible for a specific feature. Algorithms could be merged, intersected or excluded to change their behaviours. And tools such as brushes could intelligently edit code to refactor, add comments, or change the style. In fact, this idea was recently experimented with “Code Brushes” to enhance the readability or fix simple bugs in code (Wattenberger, 2024).
@@ 6. Conclusion
We presented Textoshop, a text editor that repurposes features from drawing software to offer new features and workflows that facilitate text editing. Results of a user study showed participants were more successful and preferred Textoshop to accomplish text editing tasks. The study also highlighted participants’ abilities to transfer their knowledge of drawing software to understand how to use Textoshop. Broadly, our work highlights the potential of software analogies to inform the design of software from any domain.
@@ Acknowledgements.
This work was supported by NAVER corporation as a part of NAVER-Wattpad-University of Toronto research center and NSERC under grant RGPIN-2018-05072.

### v2

@@ 5.1.3. Identifying changes was sometimes challenging with Textoshop
Some participants noted how a main difference with drawing software was that it is not as immediate to perceive textual changes. P8 commented “in like Figma […] it’s immediately, visually obvious what changed. […] It didn’t feel like I had the same mobility here where I could like immediately understand what the changes and the consequences were”. In fact, some participants mentioned how they would prefer to have some time to review the changes before integrating them. P8 added “I really want to see things side-by-side”. P6 explained that they would like a system to track the changes because the current animation of changes “disappeared so quickly before I can notice [the changes]”. In Textoshop, changes can be tracked either by working on copies or using layers, but these comments seem to indicate that, unlike drawing software, this is not enough to help users identify textual changes.
@@ 5.2.1. Textoshop might not be easy to learn for people with no experience with drawing software
5.2.2. Due to the implementation of Textoshop, it inherits all the issues of using LLMs to edit text
While the implementation of Textoshop is not core to our contribution, it is important to acknowledge that Textoshop currently heavily relies on OpenAI’s GPT-4o. It is known that LLMs have issues in terms of biases, privacy, and plagiarism (Weidinger et al., 2021; Meyer et al., 2023). We also found the model to perform poorly on tasks such as resizing to a specific number of words. In our implementation, we workaround these issues by regenerating multiple versions until one that matches the length is found. But it is often imperfect, which might frustrate users. Finally, some text transformations might result in imperfect, incongruent, or hallucinated text. These issues might be mitigated by using fine-tuned LLMs that take greater care in preserving semantic meaning and matching the style of the original text.
@@ 5.2.3. The drawing software analogy might confuse users
Using interface analogies might lead to mismatches between the source and the target domain. While some mismatches might be beneficial and lead to curiosity, others might cause problems and create invalid assumptions (Neale and Carroll, 1997). With Textoshop, this difficulty is coupled with the use of an LLM that might produce unexpected text, making the system difficult to model for the user (Chung et al., 2021). While this did not seem to be an issue during the study, some participants did mention a mismatch between their mental model and the behaviour of the system regarding the layer system. For example, P3 mentioned not seeing text layers like stacked sheets but more like “variables or conditions that you’re kind of swapping in and out”. One possible explanation is given by Neale and Carroll (1997) in that “mismatches occur with aspects of computer functionality when the target domain is unable to provide a direct real-world correlate”. Text layers have no real-world equivalent. Unlike physical transparent sheets, text layers essentially replace text and are positioned dynamically based on their anchor points in the text. Finally, text layers also led users to make invalid assumptions. For example, P8 expected features and were surprised not to find them “In Figma […] when I click again it goes down a layer rather than staying at the same level”
@@ 5.3.1. Repurposing other digital drawing features
We only scratched the surface of drawing features that could be repurposed for text. For example, layers could support different blending modes to either overwrite or track the changes and changing the opacity of the layer could control how the text is integrated (e.g., 100% overwrites, while 50% generates text with a meaning midway between the layer and what it overlaps). Other layer types could include “Adjustment layers” (Andrews, 2006) that apply a transformation to the text (e.g., making the text more positive) while retaining the integrity of the text underneath. The rotation feature could allow specifying the centre of rotation (the word that should remain at the same location). Finally, text could be transformed via histograms showing text properties such as the number of adverbs, the sentiment of the text, or the length of the sentences. Then, writers could manipulate the histogram to transform the text.
@@ 5.3.2. Leveraging recent research in graphical editing to also benefit Textoshop
The design of drawing software is a well-researched topic. Much of this research could be adapted and implemented in Textoshop. For example, work on reifying (Beaudouin-Lafon and Mackay, 2000) graphical attributes into manipulable objects (Xia et al., 2016) could be adapted so that properties such as tone, sentence length, and sentence rotation become first-class objects that can be easily manipulated, copied, reversed, and removed. Work on better colour pickers and colour theme authoring could similarly power better tone pickers where tones become objects that can be rearranged and interpolated (Shugrina et al., 2019). Similarly, work on using histograms as another editing space could inform the design of text histograms that can query and edit the text (Chevalier et al., 2012).
5.3.3. Testing the impact of Textoshop over a longer period of time
While challenging due to the nondeterministic nature of LLMs that might confound the results, a longitudinal or diary study of writers using Textoshop over several months would tell us how people integrate the features in their workflows, how they co-adapt with the system (Mackay, 2000), and how it impacts the ideation, the quality of the result, and the productivity (Chung et al., 2021). For example, the consequences of constantly switching between keyboard and pointing devices are unclear. Moving from the keyboard to the mouse takes about 0.4s (Card et al., 1978) and might have a cognitive cost which might add up in the long run. However, we note that these switches may not be so frequent considering writing is not only about typing characters on a keyboard but also involves planning and reviewing, which do not leverage the keyboard (Flower and Hayes, 1981). Additionally, Buschek (2024) observed a shift in how writing is done to look more like an assemblage of text snippets rather than manual writing. Such workflow is less reliant on the keyboard and might be better supported by Textoshop. Beyond productivity, a longitudinal study would also help study how Textoshop changes the writing workflow and the text that is produced. For example, the diverse ways participants used layers to structure their text is reminiscent of fluid documents (Chang et al., 2000) where the added structure and ways to reveal in-context information changed people’s reading behaviours (Zellweger et al., 2000). While some features of Textoshop may not make writing easier nor faster, they might contribute to other aspects such as idea generation or quality of the craft (Chung et al., 2021).
@@ 5.3.4. Applying the drawing software analogy to other domains
Analogies used to design software often stay within the same domain: the digital canvas for drawing software and the blank sheet for text editors. Through this work, we show the potential of using analogies from different domains. We believe this strategy could inform the design of software from all domains. For example, code editors could also benefit from layers to structure the code. This would allow an organization with a finer granularity than files, and layers could be hidden to remove the code responsible for a specific feature. Algorithms could be merged, intersected or excluded to change their behaviours. And tools such as brushes could perform smart editing of code to refactor, add comments, or change the coding style. In fact, this idea was recently experimented with “Code Brushes” that can enhance the readability or fix bugs in code (Wattenberger, 2024).
@@ 6. Conclusion
We presented Textoshop, a text editor that repurposes features from drawing software to offer new workflows and features that facilitate text editing. Results of a user study showed that participants were more successful and preferred Textoshop to accomplish text editing tasks. The study also highlighted how participants transferred their knowledge of drawing software to learn how to use Textoshop. Broadly, our work highlights the potential of software analogies to inform the design of software from any domain.
@@ Acknowledgements.
This work was supported by NAVER corporation as a part of NAVER-Wattpad-University of Toronto research center and NSERC Discovery Grant RGPIN-2018-05072.